OpenStudy (anonymous):

MEDAL What is the missing side length (x) Please don't tell me to use the Pythagorean theorem because I haven't learned that yet and if I haven't and they gave me this question there has to be another way to solve this

OpenStudy (anonymous):



OpenStudy (anonymous):

@k_lynn

OpenStudy (anonymous):

a^2+b^2=c^2 x^2+9^2=14^2

OpenStudy (anonymous):

x=sqrt(115)

OpenStudy (anonymous):

There doesn't seem to be a way to solve this without the pythagorean theorem or the law of sines.

OpenStudy (anonymous):

There is because I haven't learned it yet and its still asking me to answer the question

OpenStudy (anonymous):

@RainCloud @ineedurhelpnow

OpenStudy (anonymous):

The picture didn't show up right - is 14 just one of the sides, or the hypotenuse?

OpenStudy (jdoe0001):

use the pythagorean theorem, as aforementioned

OpenStudy (anonymous):

If 14 is just one of the sides and not the hypotenuse, then there's one unique Pythagorean triplet - but like everyone else has said, I don't know of a way to solve this without using the P. Theorem otherwise....

OpenStudy (anonymous):

:(

OpenStudy (triciaal):

did you learn Hero's formula for the area of a triangle given the length of the sides?

OpenStudy (triciaal):

please let me know how to do this without Pythagoren theorem, thanks

OpenStudy (anonymous):

OH! The picture of this finally showed up correctly! I didn't know that you had the value of the hypotenuse and also one of the sides. Have you done trigonometric functions in class like sine, cosine and tangent? (this might all be very much after the fact). If so, you can use them to find the missing side length.